OpenStudy (anonymous):

hello guys can some one please help me? i cant seam to solve this question... Solve the exponent equation? 5^2x=17

OpenStudy (anonymous):

You should take the natural log of both sides

OpenStudy (anonymous):

how do i do that ? can you please explain

OpenStudy (anonymous):

ok sure!

OpenStudy (anonymous):

\[\ln 5^{2x}= \ln 17\]

OpenStudy (anonymous):

\[2x \ln 5 = \ln 17 \]

OpenStudy (anonymous):

2x = (ln 17) / (ln 5)

OpenStudy (anonymous):

you should be able to solve for x from here